Give the equation, in polynomial form, of the circle if (-17,7) and (13,-9) are endpoints of a diameter of the circle.
Answer by stanbon(75887)

Give the equation, in polynomial form, of the circle if (-17,7) and (13,-9) are endpoints of a diameter of the circle.
-----------------------------
The midpoint of that diameter is (x,y)
where
x = (13-17)/2 = -2
y = (7-9)/2 = -1
-----------------------
So the center of the circle is (-2,-1)
------------------------
radius = distance from (-2,-1) to (13,-9)
= sqrt[(-9--2)^2 + (13--2)^2]
--------
= sqrt[7^2 + 15^2]
--------
= sqrt[274]
-----
Equation: (x-2)^2 + (y+1)^2 = 274
===================================
Cheers,
Stan H.
==================